# Shapes and orientations of massive halos in the statistically anisotropic universe

### v1

@@ Abstract
We investigate how statistical anisotropy (SA) in matter distributions affects the shapes and orientations of cluster-sized halos, using cosmological $N$-body simulations that incorporate SA. While the three-dimensional halo shape parameters show little dependence on SA, we find that halo orientations are significantly influenced, with halos tending to align either perpendicular or parallel to the SA direction. This SA-induced alignment becomes more prominent for more massive halos. Our findings suggest that observational measurements of projected halo shapes, such as those derived from galaxy cluster-galaxy lensing, could provide a novel probe of SA in the universe.
I introduction
@@
SA can arise from anisotropic sources such as vector fields, which could appear in anisotropic inflationary scenarios (see [1, 2, 3] for review), vector dark matter [4, 5, 6], and vector dark energy models [7]. In particular, inflationary models with gauge fields, including the anisotropic inflation model, have been extensively discussed, e.g., in the context of axion cosmology based on the string theory, primordial magnetogenesis, and so on. Thus, exploring the SA may indirectly provide a means to probe the matter contents such as gauge fields in the early universe.
@@
The quadrupolar type of SA, the leading-order term [8], is characterized by only one magnitude parameter $g_{*}$, which has been constrained by the CMB measurements of Planck satellite observations [9, 10, 11], yielding $|g_{*}|=\mathcal{O}(10^{-2})$. Additional efforts have been made to constrain SA using galaxy clustering measurements [12, 13] (see Refs. [14, 15, 16] for the practical analysis techniques using the polypolar spherical harmonic basis). Recently, Ref. [17] performed cosmological $N$-body simulations incorporating quadrupolar SA, and showed that SA induces anisotropic halo bias in the quadrupole moments of galaxy two-point statistics. Including this newly identified effect in the analysis enables more accurate constraints on $g_{*}$ beyond linear theory from galaxy clustering.
@@
Since SA introduces a preferred direction in the initial conditions of matter distributions, and halo formation traces the underlying matter distribution, it is reasonable to expect that halo shapes and orientations may be altered and become somewhat anisotropic in the statistically anisotropic universe. Observationally, the projected shapes of massive halos have been studied mainly via galaxy cluster-galaxy lensing. For example, Ref. [18] measured the distribution of the projected ellipticity of massive halos using galaxy cluster-galaxy lensing signals and found consistency with predictions from cosmological simulations by Ref. [19]. Ref. [20] obtained the mean ellipticity of galaxy clusters using distributions of member galaxies as well as weak lensing. If SA affects halo shapes and/or orientations, the observable quantities, such as the projected halo ellipticity, would also exhibit systematic directional features, which could be detected through gravitational lensing observations. In such a case, comparisons between predictions from cosmological simulations with SA and observations of projected ellipticities could serve as a means to constrain $g_{*}$.
While Ref. [17] focused on SA-induced halo bias, in this work, we investigate how SA affects the shapes and orientations of cluster-sized halos at $z=0$ using cosmological simulations [19, 21, 22], with the aim of exploring their potential as a new probe of SA. We find that the three-dimensional shape parameters [21] show surprisingly little dependence on the quadrupolar SA parameter $g_{*}$. In contrast, halo orientations are more strongly influenced by SA. Specifically, halos tend to increasingly align perpendicular to the SA-preferred direction for positively larger $g_{*}$, and parallel for negatively larger $g_{*}$. We also find that this alignment effect becomes more pronounced for more massive halos. Given that halo structures encode information about the underlying matter field, studying their response to SA offers a complementary approach to existing clustering-based analyses. This motivates our investigation of halo shapes and orientations as potential observational tracers of SA.
@@
The rest of this paper is organized as follows. In Sec. II, we describe the cosmological $N$-body simulations incorporating SA, the halo catalogs used in this work, and the measured quantities related to the halo shapes. Sec. III presents our results on the halo shapes and orientations. We summarize and conclude in Sec. IV.
II setup
In this section, we briefly describe how to quantify the SA and how to obtain the halo catalog in the simulated universe with the SA. Then, we introduce the quantities measured in simulations that characterize the shape and orientation of identified halos.
II.1 Cosmological $N$-body simulations with SA
@@
We follow Ref. [17] to perform the cosmological $N$-body simulations with SA. First let us introduce the SA in the matter distributions. In our study, we consider the quadrupolar type of SA, and the power spectrum of initial linear matter overdensity fields is written by $\left\langle\delta_{\mathrm{m}}(\mathbf{k}_{1})\delta_{\mathrm{m}}(\mathbf{k}_{2})\right\rangle=(2\pi)^{3}\delta^{(3)}(\mathbf{k}_{1}+\mathbf{k}_{2})\mathit{P}_{\mathrm{m}}(\mathbf{k}_{1})$ with
$\mathit{P}_{\mathrm{m}}(\mathbf{k})=\left[1+\frac{2}{3}g_{*}\mathcal{L}_{2}(\mu)\right]\bar{P}_{\mathrm{m}}(k),$ (1)
@@
where $g_{*}$ descibes the magnitude of SA, $\bar{P}_{\mathrm{m}}(k)$ means the isotoropic component of the matter power spectrum, $\mathcal{L}_{2}\equiv\frac{1}{2}(3\mu^{2}-1)$ with $\mu\equiv\hat{\mathbf{k}}\cdot\hat{\mathbf{d}}$ and $\hat{\mathbf{k}}\equiv\mathbf{k}/|\mathbf{k}|$, and $\hat{\mathbf{d}}$ is the directions related to SA.
@@
SA in matter distributions is incorporated by using Eq. (1) when generating the initial conditions (ICs), where the direction of SA is set to $\hat{\mathbf{d}}=(0,~{}0,~{}1)$. We use the Boltzmann code CAMB [23] to compute the isotropic linear matter power spectrum $\bar{P}_{\mathrm{m}}(k)$ in Eq. (1) at the initial redshift $z_{\rm ini}=31$. Based on the computed $\bar{P}_{\mathrm{m}}(k)$, ICs are generated using the second-order Lagrangian perturbation theory [24, 25, 26]. We input the ICs into the cosmological $N$-body solver Gadget-2 [27, 28] to follow the evolution of matter distributions in statistically anisotropic universes.
@@
Table 1: Summary of the simulation settings: $N_{\rm part}$, $L_{\rm box}$, $m_{\rm part}$, and $g_{*}$ stand for the number of simulation particles, the size of the simulation box, the mass of a simulation particle, and the magnitude of the quadrupolar-type SA, respevtively
. run $N_{\rm part}$ $L_{\rm box}$ $[h^{-1}{\rm Gpc}]$ $m_{\rm part}$ $[h^{-1}M_{\odot}]$ $g_{*}$ L05 $512^{3}$ $0.5$ $8.16\times 10^{10}$ $0,~{}\pm 0.1,~{}\pm 1$ L2 $1024^{3}$ $2$ $6.53\times 10^{11}$ $0,~{}\pm 0.1,~{}\pm 0.5,~{}\pm 1$ L4 $1024^{3}$ $4$ $5.22\times 10^{12}$ $0,~{}\pm 0.1,~{}\pm 1$
@@
Except for assuming Eq. (1) as the initial matter power spectrum, in the $N$-body simulation, we assume the standard flat-$\Lambda$ cold dark matter cosmology with the Planck’s best fit cosmlogical parameters as $\Omega_{\rm m0}=0.3156,~{}\Omega_{\Lambda 0}=0.6844,~{}H_{0}=100h=67.27~{}{\rm km~{}s^{-1}~{}{\rm Mpc}^{-1}},~{}n_{\rm s}=0.9645$ and $A_{\rm s}=2.2065\times 10^{-9}$ [29]. We perform three simulations characterized by the number of particles $N_{\rm part}$ and the box size $L_{\rm box}$, as summarized in Table 1. To study the $g_{*}$-dependence of halo properties, for each simulation we vary $g_{*}$ values as listed in the table. For each simulation setting, we conduct four realizations, resulting in a total of 68 realizations.
@@
Figure 1: Visualizations of the matter distributions at the redshift $z=0$ in the $x$-$z$ plane, in the simulated universes with $g_{*}=+1$ (left), 00 (middle), and $-1$ (right). Each panel shows a $50~{}h^{-1}{\rm Mpc}$-thick slice along the $y$-direction taken from the realization with the same initial random seed in the L05 run. The SA direction $\hat{\mathbf{d}}$ is shown in each panel.
@@
To facilitate a visual understanding of the impact of $g_{*}$, Fig. 1 shows the matter distributions at $z=0$ in the $x$-$z$ plane for the simulated universes with $g_{*}=+1$ (left), 00 (middle), and $-1$ (right). Each panel describes a $50~{}h^{-1}{\rm Mpc}$ thick slice along the $y$-direction taken from the realization with the same initial random seed in the L05 run. Compared to the middle panel with $g_{*}=0$ (i.e., the isotropic case), the matter distribution is elongated perpendicular to the SA direction $\hat{\mathbf{d}}$ for $g_{*}=+1$ (left), and along $\hat{\mathbf{d}}$ for $g_{*}=-1$ (right).
II.2 Halos as ellipsoids
@@
To identify halos in the simulated matter distributions, we use the halo finder code Rockstar [30]. As the halo mass, we employ the mass within a region whose average density is 200 times the cosmological matter density, denoted as $M_{\rm 200b}$. Since measuring halo shapes requires a sufficient number of particles, we analyze only halos that contain more than approximately 200 particles. On the other hand, a larger simulation box is needed to obtain more massive halos due to their low abundance. Taking these factors into account, we use halos with masses of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~{}13.5]$$(14,~{}14.5]$, and $(15,~{}15.5]$ from the L05, L2, and L4 runs, respectively. [^1] Our analyses do not include subhalos.
@@
To describe the shapes of halos, we regard the halos as ellipsoids characterized by three axes, $a$, $b$, and $c$ with $a\geq b\geq c$. Based on these three axes, we introduce normalized variables characterizing the halo shape [31, 21]
$s:=\frac{c}{a},~{}q:=\frac{b}{a},~{}T:=\frac{a^{2}-b^{2}}{a^{2}-c^{2}}=\frac{1-q^{2}}{1-s^{2}},$ (2)
@@
where $s$ and $q$ are the smallest- and intermediate-to-largest axis ratios, respectively, and $T$ is the so-called triaxiality of halos. For these shape parameters, we directly use the values of $s$ and $q$ produced by Rockstar, which are measured using the weighted inertia tensor of halos [21].
@@
As for the parameter characterizing the orientations of halos, we utilize the vector assigned for each halo in the Rockstar halo catalog, $\mathbf{A}=\left(A_{x},~{}A_{y},~{}A_{z}\right)$, which corresponds to the largest axis of the ellipsoid. To clearly show the $g_{*}$-dependence of the orientations of halos, we use the arranged vector with respect to the $z$-direction (i.e., the SA direction in this work) as
$\mathbf{\hat{A}}=\left(\hat{A}_{x},~{}\hat{A}_{y},~{}\hat{A}_{z}\right)$ (3)
=$=\begin{cases}\left(A_{x},~{}A_{y},~{}A_{z}\right)/|\mathbf{A}|&\text{for $A_{z}\geq 0$}\\ \left(-A_{x},~{}-A_{y},~{}-A_{z}\right)/|\mathbf{A}|&\text{for $A_{z}<0$},\end{cases}$
@@
so that the probability distribution functions (PDFs) of the orientations of halos is uniformly random in the isotropic universe. The normalization is performed to enable comparison of results across different halo mass ranges.
III results
@@
III.1 The halo shapes
Figure 2: PDF of $s$, $P(s)$, for the three halo mass ranges of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~{}13.5],~{}(14,~{}14.5]$ and $(15,~{}15.5]$ obtained from the L05, L2, and L4 runs, respectively. Each color line shows the result for $g_{*}=0$ (blue), ±$\pm 0.1$ (green) and ±$\pm 1.0$ (orange), and the dashed (solid) lines are for the cases with the positive (negative) sign. We note that all the five curves almost overlap.
Figure 3: Same as Fig. 2 but for the PDF of $T$. The vertical dashed line in each panel indicates the boundaries between halo shape categories: “oblate” for $T<1/3$, “triaxial” for 1/$31/3<T<2/$ 3, and “prolate” for $12/3<T<$ 1. Similarly to the Fig. 2, all the five curves almost overlap.
We investigate the effect of SA on the shapes of halos, and in particular, we focus on the parameters $s$ and $T$, which respectively represent the smallest-to-largest axis ratio and triaxiality of the halo as an ellipsoid. Fig. 2 shows the $g_{*}$-dependence of the PDFs of $s$, $P(s)$, for the three halo mass ranges of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~{}13.5],~{}(14,~{}14.5]$ and $(15,~{}15.5]$ obtained from the L05, L2, and L4 runs, respectively. Each color line in this figure shows the result for $g_{*}=0$ (blue), ±$\pm 0.1$ (green) and ±$\pm 1.0$ (orange) (dashed line for the positive sign, and solid line for the negative sign). The error bars represent the standard error across the four realizations for each value of $g_{*}$ within a given run (i.e., fixed $L_{\rm box}$ and $N_{\rm part}$). We note that all the PDFs in this paper are normalized to integrate into one over the entire domain. Though we are focusing on the relatively massive halos, the shape of the PDF itself seems to be consistent with the previous studies in the statistically isotropic universe (see, e.g., Ref. [21]), e.g., the peak of $P(s)$ is around $s=0.4$, and a weak mass dependence of the peak position is also visible in this figure. Then, as for the $g_{*}$-dependence, from this figure, even in the statistically anisotropic universe with the $g_{*}$ values much larger than those currently favored by the CMB observations, it is hard to find the difference from the isotropic one. We also observe a similarly negligible deviation from the isotropic case in the PDFs of the triaxiality, $T$, $P(T)$, shown in Fig. 3. Thus, we conclude that for the cluster-sized massive halos the SA effect on the shape is negligible. We expect that the above result on the shapes can be understood as follows. As discussed in Refs. [16, 17], the power spectrum of initial linear matter overdensity fields given by Eq. (1) can be rewritten as
$\mathit{P}_{\mathrm{m}}(\mathbf{k})=\left[1+\mathcal{G}_{ij}\hat{k}_{i}\hat{k}_{j}\right]\bar{P}_{\mathrm{m}}(k)~{},$ (4)

### v2

@@ Abstract
We investigate how statistical anisotropy (SA) in matter distributions affects the distributions of shapes and orientations of cluster-sized halos, using cosmological $N$-body simulations that incorporate SA. While the three-dimensional halo shape parameters show little dependence on SA, we find that halo orientations are significantly influenced, with halos tending to align either perpendicular or parallel to the SA direction. This SA-induced alignment becomes more prominent for more massive halos. We also study other vector quantities associated with the dynamics of halos, such as bulk velocity and angular momentum vectors. We find that their dependences on the SA are smaller than those of the orientation vectors. Our findings suggest that observational measurements of projected halo shapes derived from galaxy cluster-galaxy lensing could provide a novel probe of SA in the universe.
flushbottom
1 Introduction
@@
SA can arise from anisotropic sources such as vector fields, which could appear in anisotropic inflationary scenarios (see Refs. [Dimastrogiovanni:2010sm, Soda:2012zm, maleknejad2013gauge] for review), vector dark matter [Hambye:2008bq, Graham:2015rva, Bastero-Gil:2018uel, 2025PhRvD.111j3520F], and vector dark energy models [BeltranJimenez:2008iye]. In particular, inflationary models with gauge fields, including the anisotropic inflation model, have been extensively discussed, e.g., in the context of axion cosmology based on the string theory, primordial magnetogenesis, and so on. Thus, exploring the SA may indirectly provide a means to probe such scenarios.
@@
The quadrupolar type of SA, the leading-order term [Ackerman:2007nb], is characterized by only one magnitude parameter $g_{*}$, which has been constrained by the CMB measurements of Planck satellite observations [2013PhRvD..88j1301K, Planck:2018jri, Planck:2019evm, Planck:2019kim], yielding $|g_{*}|=\mathcal{O}(10^{-2})$. Additional efforts have been made to constrain SA using galaxy clustering measurements [pullen2010non, sugiyama18] (see Refs. [Shiraishi:2016wec, shiraishi21, shiraishi23] for the practical analysis techniques using the polypolar spherical harmonic basis). Recently, Ref. [Masaki:2024hzn] performed cosmological $N$-body simulations incorporating quadrupolar SA, and showed that SA induces anisotropic halo bias in the quadrupole moments of galaxy cluster two-point statistics. Including this newly identified effect in the analysis enables more accurate constraints on $g_{*}$ from galaxy cluster clustering.
@@
Inspired by the previous work, we, in this paper, investigate how SA affects the shapes and orientations of cluster-sized halos at $z=0$ using cosmological simulations [jing02, 2006MNRAS.367.1781A, schneider12], with the aim of exploring their potential as a new and complementary probe of SA. Since SA introduces a preferred direction in the initial conditions of matter distributions, and halo formation traces the underlying matter distribution, it is reasonable to expect that halo shapes and orientations may be altered and become somewhat anisotropic in the statistically anisotropic universe. In fact, observationally, the projected shapes of massive halos have been studied mainly via galaxy cluster-galaxy lensing. For example, Ref. [oguri10] measured the distribution of the projected ellipticity of massive halos using galaxy cluster-galaxy lensing signals and found consistency with predictions from cosmological simulations [jing02]. Ref. [shin18] obtained the mean ellipticity of galaxy clusters using distributions of member galaxies as well as weak lensing. If SA affects halo shapes and/or orientations, the observable quantities, such as the projected halo ellipticity, would also exhibit systematic directional features, which could be detected through gravitational lensing observations. In such a case, comparisons between predictions from cosmological simulations with SA and observations of projected ellipticities could serve as a means to constrain $g_{*}$. More interestingly, recent analysis of the Dark Energy Survey year-3 shape catalog [2025arXiv251110005D] has reported a coherent large-scale axial intrinsic alignment signal, which could potentially intersect with SA-related phenomena, raising the possibility that SA might contribute to or be probed through such large-scale alignment signals.
By carefully investigating the halo catalogue obtained through the cosmological $N$-body simulations with the SA, we find that the three-dimensional shape parameters [2006MNRAS.367.1781A] show little dependence on the quadrupolar SA parameter $g_{*}$. In contrast, halo orientations are more strongly influenced by SA. Specifically, halos tend to increasingly align perpendicular to the SA-preferred direction for positively larger $g_{*}$, and parallel for negatively larger $g_{*}$. We also find that this alignment effect becomes more pronounced for more massive halos. In addition to shapes and orientations, we examine the SA effect on vector quantities associated with the dynamics of halos, such as bulk velocities and angular momenta since the SA-distorted matter distributions can affect the bulk velocity fields. By comparing the SA effects on these halo orientations and dynamical vectors, we conclude that the orientations are more sensitive and serve as a possible new probe of the SA.
@@
The rest of this paper is organized as follows. In Sec. 2, we describe the cosmological $N$-body simulations incorporating SA, the halo catalogs used in this work, and the measured quantities related to the halo shapes. Sec. 3 presents our results on halo shapes, orientations, and other dynamical vector quantities such as bulk velocities and angular momenta. We summarize and conclude in Sec. 4.
2 Simulations
We first outline the formulations used to incorporate the SA into cosmological $N$-body simulations. We then describe the simulation settings and define the ellipsoidal parameters that characterize the shapes of halos identified in the SA simulations.
2.1 Cosmological $N$-body simulations with SA
@@
We follow Ref. [Masaki:2024hzn] to perform cosmological $N$-body simulations with SA. First, let us introduce the SA in matter distributions. In our study, we consider the quadrupolar type of SA. The power spectrum of initial linear matter overdensity field is written by $\left\langle\delta_{\mathrm{m}}(\mathbf{k}_{1})\delta_{\mathrm{m}}(\mathbf{k}_{2})\right\rangle=(2\pi)^{3}\delta^{(3)}(\mathbf{k}_{1}+\mathbf{k}_{2})\mathit{P}_{\mathrm{m}}(\mathbf{k}_{1})$ with
$\mathit{P}_{\mathrm{m}}(\mathbf{k})=\left[1+\frac{2}{3}g_{*}\mathcal{L}_{2}(\mu)\right]\bar{P}_{\mathrm{m}}(k),$ (2.1)
@@
where $g_{*}$ describes the magnitude of SA, $\bar{P}_{\mathrm{m}}(k)$ means the isotoropic component of the matter power spectrum, $\mathcal{L}_{2}:=\frac{1}{2}(3\mu^{2}-1)$ is the second-order Legendre polynomial with $\mu:=\hat{\mathbf{k}}\cdot\hat{\mathbf{d}}$ and $\hat{\mathbf{k}}:=\mathbf{k}/|\mathbf{k}|$, and $\hat{\mathbf{d}}$ is the directions related to SA.
@@
SA in matter distributions is incorporated using Eq. (2.1) when generating initial conditions (ICs), where the direction of SA is set to $\hat{\mathbf{d}}=(0,~0,~1)$. We use the Boltzmann code CAMB [camb] to compute the isotropic linear matter power spectrum $\bar{P}_{\mathrm{m}}(k)$ in Eq. (2.1) at the initial redshift $z_{\rm ini}=31$. Based on the calculated $\bar{P}_{\mathrm{m}}(k)$, ICs are generated using the second-order Lagrangian perturbation theory [scoccimarro98, Crocce06a, nishimichi09]. We input the ICs into the cosmological $N$-body solver Gadget-2 [gadget, gadget2] to follow the evolution of matter distributions in statistically anisotropic universes.
@@
Table 1: Summary of the simulation settings: $N_{\rm part}$, $L_{\rm box}$, $m_{\rm part}$, and $g_{*}$ stand for the number of simulation particles, the size of the simulation box, the mass of a simulation particle, and the magnitude of the quadrupolar-type SA, respectively.
run $N_{\rm part}$
$L_{\rm box}$
$h^{-1}{\rm Gpc}]$
$m_{\rm part}$
$h^{-1}M_{\odot}]$
$g_{*}$
L05 $512^{3}$ $0.5$ $8.16\times 10^{10}$ $0,~\pm 0.1,~\pm 1$
L2 $1024^{3}$ 22 $6.53\times 10^{11}$ $0,~\pm 0.1,~\pm 0.5,~\pm 1$
L4 $1024^{3}$ 44 $5.22\times 10^{12}$ $0,~\pm 0.1,~\pm 1$
@@
Except for assuming Eq. (2.1) as the initial matter power spectrum, in the $N$-body simulation, we assume the standard flat-$\Lambda$ cold dark matter cosmology with the Planck’s best fit cosmological parameters as $\Omega_{\rm m0}=0.3156,~\Omega_{\Lambda 0}=0.6844,~H_{0}=100h=67.27~{\rm km~s^{-1}~{\rm Mpc}^{-1}},~n_{\rm s}=0.9645$ and $A_{\rm s}=2.2065\times 10^{-9}$ [planck-collaboration:2015fj]. We perform three simulations characterized by the number of particles $N_{\rm part}$ and the box size $L_{\rm box}$, as summarized in Table 1. To study the $g_{*}$-dependence of the halo properties, for each simulation we vary the $g_{*}$ values as listed in the table. For each simulation setting, we conducted three realizations, resulting in a total of 51 realizations.
@@
Figure 1: Visualizations of the matter distributions at the redshift $z=0$ in the $x$-$z$ plane, in the simulated universes with $g_{*}=+1$ (left), 0 (middle), and 1-1 (right). Each panel shows a $50~h^{-1}{\rm Mpc}$-thick slice along the $y$-direction taken from the realization with the same initial random seed in the L05 run. The SA direction $\hat{\mathbf{d}}$ is shown in each panel.
@@
To facilitate a visual understanding of the impact of $g_{*}$, Fig. 1 shows the matter distributions at $z=0$ in the $x$-$z$ plane for the simulated universes with $g_{*}=+1$ (left), 0 (middle), and 1-1 (right). Each panel describes a $50~h^{-1}{\rm Mpc}$ thick slice along the $y$-direction taken from the realization with the same initial random seed in the L05 run. Compared to the middle panel with $g_{*}=0$ (i.e., the isotropic case), the matter distribution is elongated perpendicular to the SA direction $\hat{\mathbf{d}}$ for $g_{*}=+1$ (left) and along $\hat{\mathbf{d}}$ for $g_{*}=-1$ (right).
2.2 Halos as ellipsoids
@@
To identify halos in the simulated matter distributions, we use the halo finder code Rockstar [Behroozi:2013]. We define the halo mass as the mass enclosed within a region whose mean density is 200 times the background matter density, denoted as $M_{\rm 200b}$. Since measuring halo shapes requires a sufficient number of particles, we analyze only halos that contain more than approximately 200 particles. On the other hand, a larger simulation box is required to obtain more massive halos due to their lower abundance. Taking these factors into account, we use halos with masses of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~13.5]$(14,14.5](14,~14.5], and (15,15.5](15,~15.5] from the L05, L2, and L4 runs, respectively. [^1] Our analyses do not include subhalos.
@@
To describe the shapes of halos, we regard the halos as ellipsoids characterized by three axes, $a$, $b$, and $c$ with $a\geq b\geq c$. Based on these three axes, we introduce normalized variables that characterize the halo shapes [franx91, 2006MNRAS.367.1781A]
$s:=\frac{c}{a},~q:=\frac{b}{a},~T:=\frac{a^{2}-b^{2}}{a^{2}-c^{2}}=\frac{1-q^{2}}{1-s^{2}},$ (2.2)
@@
where $s$ and $q$ are the smallest- and intermediate-to-largest axis ratios, respectively, and $T$ is the so-called triaxiality of halos. For these shape parameters, we directly use the values of $s$ and $q$ produced by Rockstar, which are measured using the weighted inertia tensor of halos [2006MNRAS.367.1781A].
@@
As for the parameter characterizing the orientations of halos, we utilize the vector assigned for each halo in the Rockstar halo catalog, $\mathbf{A}=\left(A_{x},~A_{y},~A_{z}\right)$, which corresponds to the orientation of the largest axis of the ellipsoid times the halo size. To clearly show the $g_{*}$-dependence of the orientations of halos, we use the arranged vector with respect to the $z$-direction (i.e., the SA direction in this work) as
$\mathbf{\hat{A}}=\left(\hat{A}_{x},~\hat{A}_{y},~\hat{A}_{z}\right)$ (2.3)
[…]
@@
so that the probability distribution functions (PDFs) of the orientations of halos are uniformly random in the isotropic universe. The normalization is performed to enable comparison of results across different halo mass ranges.
3 Results
@@
3.1 Halo shapes
Figure 2: Left: PDFs of $s$, $P(s)$, for the three halo mass ranges of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~13.5],~(14,~14.5]$ and (15,15.5](15,~15.5] obtained from the L05, L2, and L4 runs, respectively. Each color line shows the result for $g_{*}=0$ (blue), $\pm 0.1$ (green), and $\pm 1.0$ (orange), and the dashed (solid) lines are for the cases with the positive (negative) sign. We note that all five curves almost overlap. Right: Same as the left panel but for the PDFs of $T$, $P(T)$. The vertical dashed line in each panel indicates the boundaries between halo shape categories: “oblate” for $T<1/3$, “triaxial” for $1/3<T<2/3$, and “prolate” for $2/3<T<1$. Similar to the left panel, all five curves almost overlap.
We investigate the effect of SA on the shapes of halos, and in particular, we focus on the parameters $s$ and $T$.
The left panel of Fig. 2 shows the $g_{*}$-dependence of the PDFs of $s$, $P(s)$, for the three halo mass ranges of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~13.5],~(14,~14.5]$ and (15,15.5](15,~15.5] obtained from the L05, L2, and L4 runs, respectively. Each color line in this figure shows the result for $g_{*}=0$ (blue), $\pm 0.1$ (green), and $\pm 1.0$ (orange), with dashed and solid lines indicating the positive and negative signs, respectively. The error bars represent the standard error across the three realizations for each value of $g_{*}$ within a given run (i.e., fixed $L_{\rm box}$ and $N_{\rm part}$). We note that all the PDFs in this paper are normalized to integrate into one over the entire domain. Though we focus on the relatively massive halos, the shape of the PDF itself seems to be consistent with the previous studies in the statistically isotropic universe (see, e.g., Ref. [2006MNRAS.367.1781A]), e.g., the peak of $P(s)$ is around $s=0.4$, and a weak mass dependence of the peak position is also visible in this figure.
As for the $g_{*}$-dependence, from this figure, even in the statistically anisotropic universe with the $g_{*}$ values much larger than those currently favored by the CMB observations, it is hard to find the difference from the isotropic one. We also observe a similarly negligible deviation from the isotropic case in the PDFs of the triaxiality, $P(T)$, shown in the right panel of Fig. 2. Thus, we conclude that for the cluster-sized massive halos, the SA effect on the shape is negligible. We expect that the above results on the shapes can be understood as follows. As discussed in Refs. [shiraishi23, Masaki:2024hzn], the power spectrum of initial linear matter overdensity fields given by Eq. (2.1) can be rewritten as
$\mathit{P}_{\mathrm{m}}(\mathbf{k})=\left[1+\mathcal{G}_{ij}\hat{k}_{i}\hat{k}_{j}\right]\bar{P}_{\mathrm{m}}(k)~,$ (3.1)